A researcher randomly picked 50 students from a small University and found the followings;

Average GPA = 2.8

Variance = 0.04

  1. Given the information above the researcher calculated that 68% of the 50 students have a GPA between ____________ and ___________.

2. Using the information above, if Lisa has a GPA of 3.5, calculate the Z score for Lisa.
Answer #1

Solution:

Given, the Normal distribution with,

   = 2.8

2 = 0.04

So , = 0.04 = 0.2

1) According to the Empirical rule ,

68% observations lie between - and +

95% observations lie between - 2 and + 2

99.7% observations lie between - 3 and + 3

- = 2.8 - 0.2 = 2.6

+ = 2.8 + 0.2= 3.0

Answer : 68% of the 50 students have a GPA between 2.6 and 3.0

2)

Lisa has a GPA of 3.5

X = 3.5

Now ,

Z score = (X - )/ = (3.5 - 2.8)/0.2 = 3.50

z score for Lisa is 3.50